# Application of Transformers for Nonlinear Channel Compensation in Optical Systems

### v2

@@ Abstract
In this paper, we introduce a new nonlinear optical channel equalizer based on Transformers. By leveraging parallel computation and attending directly to the memory across a sequence of symbols, we show that Transformers can be used effectively for nonlinear equalization in coherent long-haul transmission. For this application, we present an implementation of the encoder part of the Transformer and analyze its performance over a wide range of different hyper-parameters. It is shown that by processing blocks of symbols at each iteration and carefully selecting subsets of the encoder’s output to be processed together, an efficient nonlinear compensation can be achieved for different complexity constraints. We also propose the use of a physic-informed mask inspired by nonlinear perturbation theory for reducing the computational complexity of the attention mechanism.
1 Introduction
@@
The impact of nonlinear interference from the fiber Kerr effect and optical/electrical components is a main challenge for high-speed optical systems that limit the achievable transmission rates. In general, fiber nonlinearity interference can be equalized in sample or symbol domain by approximating and inversing the nonlinear Schrodinger equation through deterministic models such as digital back-propagation (DBP) [1, 2, 3] or perturbation-based nonlinear compensation (PNLC) methods [4, 5]. Alternatively, various solutions have been proposed in order to exploit deep artificial neural networks (ANN) for nonlinear compensation in optical communications [6, 7, 8]. The pattern and medium dependent characteristics of nonlinear propagation make it a suitable case for artificial neural network domain. Specifically, recurrent neural networks (RNN) such as long short-term memory (LSTM) operating on received symbol sequences have shown to be effective in fiber nonlinearity compensation for different transmission scenarios. [9, 10, 11]. Attempts are also made to design appropriate ANNs for hardware implementation with real-world applications in mind [12, 13]. However, the sequential nature of RNNs at the core of better performing architectures, is a bottleneck for parallel implantation that is vital for hardware developments of ultra-high-speed optical transmission systems.
@@
On the other hand, Transformers as powerful ANN structures based on self-attention mechanism have been introduced in [14]. Since their birth, these structures are widely employed in various machine learning applications with impressive performance. In fact, Transformers are designed to overcome the limits of sequential nature of RNNs. In particular, unlike RNNs, Transformers are highly parallelizable in both training and inference stage which makes them greatly suitable for hardware developments of ultra-high-speed optical transmissions. In this work, we employ Transformers for applications in coherent optical communications. To the best of our knowledge for the first time, we study the application of Transformer as the core of fiber nonlinear equalization and explore various opportunities and challenges associated to this architecture. We show that in order to leverage parallel computation and attending directly to the memory across a sequence of received symbols, Transformers with efficient embeddings and implementation can be used for nonlinear compensation. One major advantage of Transformers over LSTMs for nonlinear compensation (NLC) application is that both the training and inference is performed over a block of received symbols. RNNs generally suffer from the vanishing gradient during the training stage as well as increased latency in the inference stage over long blocks of symbols [12]. In contrast, Transformer-NLC can be directly optimized over the target block length and is inherently more suitable for applications in high-speed coherent optical transceivers that require a great deal of parallelization to satisfy their high-throughput and low-latency requirements. Furthermore, we present the use of a physic-informed mask according to the nonlinear perturbation theory in order to make the attention matrix sparse. A two-dimensional relations modeled as a mask matrix highlights the importance of combinations of symbols for the nonlinear estimation and provides savings in the computational complexity specially at regions with very limited resources.
@@
The remainder of this paper is organized as follows: In Section 2, the overall structure of fiber nonlinear compensation is briefly discussed and the Transformer’s basics are reviewed. In Section 3, the NLC architecture with Transformers are introduced for processing a block of symbols. Moreover, the impact of the perturbation mask for reducing the computational complexity of the attention mechanism is explained. Next, numerical results are presented in Section 4 and finally, we summarize the paper in Section 5.
2 Preliminaries
2.1 Nonlinear Compensation in Coherent Optical Systems
@@
where $u_{x/y}=u_{x/y}(t,z)$ represents the optical field of polarization $x$ and $y$, respectively, $\alpha$ is the attenuation coefficient, $\beta$ is the group velocity dispersion (GVD), and $\gamma$ is the nonlinear coefficient. Kerr nonlinear impact can traditionally be equalized through deterministic nonlinear models by approximating and inversing the Eq. (1) through DBP [1, 2, 3] where the fiber is modeled as a series of linear and nonlinear sections through first-order approximation of Manakov equation. Also, by employing the perturbation analysis [4], the optical field can be represented as the solution of a linear term plus a nonlinear perturbation term in symbol domain and in a lumped stage. In fact, the first-order perturbation term can be modeled as the weighted sum of triplets of symbols in addition to a constant phase rotation[5, 16].
Figure 1: Receiver symbol-domain diagram for neural network nonlinear compensation.
Various solutions are introduced to employ machine learning for the fiber nonlinearity compensation(e.g. [6, 7, 8, 9, 10, 11]) in which one can learn the equalization process through data. These artificial neural network nonlinear compensation (ANN-NLC) solutions include learned-DBP, using triplets of symbols for PNLC with basic neural networks, and using soft-symbols in deep neural networks. Furthermore, a data-driven ANN solution provides a more general and flexible approach while giving us the potentials for large simplifications in computational complexity with efficient quantization of the network [17].
A basic block diagram for a receiver side equalization process is presented in Fig. 1 where after the conventional linear processing for the coherent optical receiver, the preprocessing buffer generates appropriate inputs for an ANN method. This can be seen in form of calculated PNLC triplets for a perturbation-based method, or a vector of soft symbols at the end of Rx-DSP chain for a deep artificial neural network nonlinear compensation. Assuming the first-order perturbation is the dominant term, a scaler can be used to adjust the model for different optical launch powers:
@@
2.2 Transformers
Transformers were introduced in [14] based on the self-attention mechanism and have been employed in different fields of machine learning including language modeling, vision, audio, and time series processing with impressive performance while being appropriate for parallel implementation. Specifically, a Transformer processes all symbols in parallel and provides direct interactions among symbols in an input sequence which makes it capable of capturing memories more efficiently compared to an LSTM where the memory is handled sequentially. This makes Transformers highly matched for real-time applications such as hardware developments of ultra-high-speed optical transmissions.
A vanilla Transformer, originally developed for natural language processing, consists of an encoder and decoder architecture. However, for our application of nonlinear equalization, we just employ the encoder part of a Transformer which is depicted in Fig. 2. This architecture consists of a module that generates the embeddings, a positional encoder, $L$ stacked layers of multi-head attention (MHA) and position-wise feed-forward network (FFN), and an output module at end that can be an MLP to generate the estimated nonlinear interference associated to real and imaginary parts of the equalized output symbol sequence. Furthermore, there are residual connections and a layer normalization at each Transformer layer.
Figure 2: Transformer encoder architecture.
@@ 2.2 Transformers
Transformers also use residual connections and a layer normalization after each pair of attention and FNN modules as described by:
@@
3 Transformers for Nonlinear Compensation
3.1 Model Architecture
In this work, we use the above Transformer encoder architecture to estimate the nonlinear distortions in the received symbols. The overall design for the nonlinear channel equalization is depicted in Fig. 3. Due to the channel memory and pattern-dependency over the transmitted sequence, it is required to process several symbols before and after each target symbol. The number of these extra symbols from each side of the target symbol is denoted by tap-size value $t$. The model uses received symbols from four components of the optical domain at the end of the DSP chain, $X_{I},X_{Q},Y_{I},Y_{Q}$. At first, an embedding generator module maps this sequence of received symbols into $R^{N\times d_{model}}$, where $N$ is the embedding sequence length. Next, a Transformer is used as the core for channel nonlinear equalization. The dashed block in Fig. 2 corresponds to the Transformer block that operates on the generated embeddings.
Figure 3: The overall model architecture for the nonlinear channel equalization. From left to right: the CNN generates input embeddings which are processed by the Transformer to generate output representations. These outputs are then fed into an MLP to generate the estimated nonlinear distortions $E_{XI}$ and $E_{XQ}$.
The next module processes the Transformer’s output to generate the equalizer’s nonlinear distortion estimations. By using a Transformer, we can process $N$ input embeddings simultaneously and produce $N$ output representations. However, since we are interested in computing the distortion for a target symbol $s$, we may not need all the $N$ output representations and a subset of them may suffice. In other words, in order to produce the nonlinear distortion for symbol $s$, we may only use the Transformer’s output representation for $s$ or use a window of representations around $s$. We call these extra representations the neighbors of $s$. The impact of optimizing this neighboring window of size $W$ on the design is explained more in Section 3.4. Specifically, the selected window of representations are fed into an MLP module consists of a linear layer ($\in R^{W\cdot d_{model}\times 2}$), a leaky ReLU nonlinear activation function with negative slope of $0.2$, a second linear layer ($\in R^{2\times 10}$), a leaky ReLU nonlinear activation function with negative slope of $0.2$, and a third linear layer ($\in R^{10\times 2}$). The last layer has two outputs that generates $E_{XI}$ and $E_{XQ}$, the estimated nonlinear distortions corresponding to the target symbol.
@@ 3.1 Model Architecture
For processing a block of symbols for equalization, the training/inference samples are created by selecting a block of $b$ symbols. In order to maintain the quality of equalization at the beginning and end of the block, we append extra symbols to the beginning and end of the block; effectively expanding the selected symbols with extra $2t$ symbols at each side. Hence, each input sequence length is equal to $b+2t$. This sequence of symbols passes through the embedding generator module which produces $b+\ell$ embeddings (the exact value of $\ell$ will be explained later). These embeddings are then fed into the Transformer where $N=b+\ell$ outputs are generated. The output module (MLP in the Fig. 3) generates $b$ estimated nonlinear distortions which are used for computing the loss in the training stage. This loss is generated as the mean of squared differences between $b$ estimated nonlinear distortions and $b$ target ones. Through back-propagation, the optimizer uses this loss to modify the weights in the training stage. With this approach, training can be accelerated by a large factor since instead of training over one symbol at a time, the model is trained on $b$ symbols at each training iteration. In addition, in the inference (equalization) stage, the Transformer generates $b$ nonlinear distortions at once through operations that are highly parallelizable. This can be seen from the following equation where the computational complexity of the attention matrix is expressed as a function of block size $b$:
$RMPS_{att-blk}=\frac{h(b+\ell)^{2}d_{k}+3h(b+\ell)^{2}}{b}.$ (6)
@@
where $h$ is the number of heads and RMPS stands for real multiplications per symbol. In analyzing the computational complexity, we only consider the multiplication operations and do not include the additions. Hence, RMPS is used as the complexity metric in this work. The first term in the above equation is the complexity of computing $QK^{T}$, and the second term is the complexity of division by $\sqrt{d_{k}}$ and softmax computation. In a vanilla Transformer, the size of $d_{K}$ and $d_{V}$ are chosen as $d_{model}/h$. In our studies, we also explore the configurations where $d_{K}$ and $d_{V}$ are equal to $d/h$ where $d$ may not be equal to $d_{model}$. This case is particularly of interest since for $d<d_{model}$, we can reduce the computational complexity. Note that, ReLU and leaky ReLU’s complexities are excluded in our analysis, since their hardware implementations are fairly simple and requires very little resources. Note that the reduction factor of $b$ also exists in the other layers including the FFNs and output layer. Assuming $\ell$ and $h$ constants, for large $b$, the overall complexity in Eq. (6) is approximately $O(bd_{k})$ per symbol which is in fact linear in block size $b$.
3.2 Embedding
@@
Our studies over the course of this research show that the Transformer which directly operates on input symbols cannot effectively learn the nonlinear interactions among the symbols. Hence, we need to create good representations, commonly known as embeddings, of the input components, $X_{I},X_{Q},Y_{I},Y_{Q}$, prior to feeding them to the Transformer. Here, we explore two main classes of ANN models for generating the embeddings, an MLP and a CNN. For the MLP architecture, we examine two configurations. In the first one, an MLP with only one linear layer is used. In the second configuration we employ a multi-layer MLP comprising a linear layer, a Leaky ReLU nonlinear activation with the negative slope of $0.2$, and another linear layer. Note that the MLPs are applied similarly to each position over the input sequence.
@@
Using these MLPs, we project the input symbols ($R^{(2t+b)\times 4}$) into embeddings with dimension $R^{(\ell+b)\times d_{model}}$ where $b$ is the processing block size, $\ell=2t$ and $t$ is the tap size as defined by the number of symbols on each side of the target symbol(s). We also explore the use of CNNs for creating embeddings since they have shown great potential for generating features from sequences. We use a CNN with one layer of one-dimensional convolutional layer with a kernel size of $k$, four input channels, $d_{model}$ output channels, and a stride of one, followed by a leaky ReLU nonlinear activation function with negative slope of $0.2$. The CNN generates embeddings of dimension $R^{(\ell+b)\times d_{model}}$ where $\ell=2t-k+1$. Note that we use $\ell$ to unify our math notation for representing computation complexities for Transformers with either MLP or CNN embeddings-generator modules. One also should note that $\ell$ is equal to $2t$ for MLP generated embeddings and $12t-k+$ 1 for CNN ones.
3.3 Mask

### v3

@@ Abstract
In this paper, we introduce a new nonlinear optical channel equalizer based on Transformers. By leveraging parallel computation and attending directly to the memory across a sequence of symbols, we show that Transformers can be used effectively for nonlinear compensation (NLC) in coherent long-haul transmission systems. For this application, we present an implementation of the encoder part of the Transformer and analyze its performance over a wide range of different hyper-parameters. It is shown that by proper embeddings and processing blocks of symbols at each iteration and also carefully selecting subsets of the encoder’s output to be processed together, an efficient nonlinear equalization can be achieved for different complexity constraints. To reduce the computational complexity of the attention mechanism, we further propose the use of a physic-informed mask inspired by nonlinear perturbation theory. We also compare the Transformer-NLC with digital back-propagation (DBP) under different transmission scenarios in order to demonstrate the flexibility and generalizability of the proposed data-driven solution.
Index Terms:
Optical Fiber Communication, Nonlinear Channel Compensation, Transformer, Attention Mechanisms
I Introduction
@@
The impact of nonlinear interference from the fiber Kerr effect and optical/electrical components is a main challenge for high-speed optical systems that limits the achievable transmission rates. In general, fiber nonlinearity interference can be equalized in sample or symbol domain by approximating and inversing the nonlinear Schrodinger equation through deterministic models such as DBP [1, 2, 3] or perturbation-based nonlinear compensation (PNLC) methods [4, 5]. Alternatively, various solutions have been proposed in order to exploit artificial neural networks (ANN) for nonlinear compensation in optical communications [6, 7, 8].
The pattern- and medium-dependent characteristics of nonlinear propagation make it a suitable case for artificial neural network domain. Specifically, recurrent neural networks (RNN) such as long short-term memory (LSTM) operating on received symbol sequences have shown to be effective in fiber nonlinearity compensation for different transmission scenarios. [9, 10, 11]. Attempts are also made to design appropriate ANNs for hardware implementation with real-world applications in mind [12, 13]. However, the sequential nature of RNNs at the core of better performing architectures, is a bottleneck for parallel implantation that is vital for hardware developments of ultra-high-speed optical transmission systems. On the other hand, Transformers as powerful ANN structures based on self-attention mechanism have been introduced in [14]. Since their birth, these structures are widely employed in various machine learning applications with impressive performance. In fact, Transformers are designed to learn long memories while overcoming the limits of sequential nature of RNNs. In particular, unlike RNNs, Transformers are highly parallelizable in both training and inference stages which makes them greatly suitable for hardware developments of ultra-high-speed optical transmissions.
@@
In this work, we employ Transformers for applications in coherent optical communications. To the best of our knowledge for the first time, we study the application of Transformer as the core of fiber nonlinear equalization and explore various opportunities and challenges associated to this architecture. We show that in order to leverage parallel computation and attending directly to the memory across a sequence of received symbols, Transformers with efficient embeddings and implementation can be used for nonlinear compensation. One major advantage of Transformers over LSTMs for nonlinear compensation application is that both the training and inference are performed over a block of received symbols. RNNs generally suffer from the vanishing gradient during the training stage as well as increased latency in the inference stage over long blocks of symbols [12]. In contrast, Transformer-NLC can be directly optimized over the target block length and is inherently more suitable for applications in high-speed coherent optical transceivers that require a great deal of parallelization to satisfy their high-throughput and low-latency requirements.
We also present the use of a physic-informed mask according to the nonlinear perturbation theory in order to make the attention matrix sparse. A two-dimensional relation, modeled as a mask matrix, highlights the importance of combinations of symbols for the nonlinear estimation and is shown to generally provide savings in the computational complexity while particularly demonstrates performance advantage at regions with limited resources. Finally, we present a comparison of performance and complexity of Transformers with DBP for nonlinear channel compensation under different transmission scenarios and demonstrate the capabilities of Transformers compared to DBP as a flexible universal solution for optical networks.
@@
The remainder of this paper is organized as follows: In Section II, the structure of fiber nonlinear compensation is briefly discussed and the Transformer’s basics are reviewed. In Section III, the Transformer-NLC architecture is introduced and then the impact of perturbation mask for reducing the computational complexity of the attention mechanism is explained. Next, numerical results are presented in Section IV and finally, we summarize the paper in Section V.
II Preliminaries
II-A Nonlinear Compensation in Coherent Optical Systems
@@
where $u_{x/y}=u_{x/y}(t,z)$ represents the optical field of polarization $x$ and $y$, respectively, $\alpha$ is the attenuation coefficient, $\beta$ is the group velocity dispersion (GVD), and $\gamma$ is the nonlinear coefficient. Kerr nonlinear impact can traditionally be equalized through deterministic nonlinear models by approximating and inversing the Eq. (1) through DBP [1, 2, 3] where the fiber is modeled as a series of linear and nonlinear sections through first-order approximation. Also, by employing the perturbation analysis [4], the optical field can be represented as the solution of a linear term plus a nonlinear perturbation term in symbol domain as a lumped stage. In fact, the first-order perturbation term can be modeled as the weighted sum of triplets of symbols in addition to a constant phase rotation[5, 16].
Figure 1: An example of receiver symbol-domain nonlinear compensation with neural network.
Applying data-driven models, various solutions are introduced to employ machine learning for the fiber nonlinearity compensation(e.g. [6, 7, 8, 9, 10, 11]) in which one can learn the equalization process through data. These artificial neural network nonlinear compensation (ANN-NLC) solutions include learned-DBP, using triplets of symbols for PNLC with basic neural networks, and using soft-symbols in deep neural networks. Basically, a data-driven ANN solution provides a more general and flexible approach while providing the potentials for large simplifications in computational complexity with efficient pruning and quantization of the network [17, 18, 19, 20].
Considering one-stage nonlinear compensation, a basic block diagram for the receiver side equalization process is presented in Fig. 1 where after the conventional linear processing for the coherent optical receiver, the preprocessing buffer generates appropriate inputs for an ANN method. This can be seen in form of calculated PNLC triplets for a perturbation-based method, or a vector of soft symbols at the end of Rx-DSP chain for a deep artificial neural network nonlinear compensation. Assuming the first-order perturbation is the dominant term, a scaler can be used to adjust the model for different optical launch powers:
@@
II-B Transformers
Transformers were introduced in [14] based on the self-attention mechanism. They have been employed in different fields of machine learning including language modeling, vision, audio, and time series processing with impressive performance while being appropriate for parallel implementation. Specifically, a Transformer processes all symbols in parallel and provides direct interactions among symbols in the input sequence which makes it capable of capturing memories more efficiently compared to an LSTM where the memory is handled sequentially. This makes Transformers highly matched for real-time applications such as hardware developments of ultra-high-speed optical transmissions.
A vanilla Transformer, originally developed for natural language processing, consists of an encoder-decoder architecture. However, for our application of nonlinear equalization, we just employ the encoder part of a Transformer which is depicted in Fig. 2. This architecture consists of a module that generates the embeddings, a positional encoder, $L$ stacked layers of multi-head attention (MHA), position-wise feed-forward network (FFN), and an output module at end that can generate the estimated nonlinear interference associated to real and imaginary parts of the equalized output symbol sequence. Furthermore, there are residual connections and a layer normalization at each Transformer layer.
@@ II-B Transformers
Transformers also use residual connections and layer normalization modules after each pair of attention and FNN modules as described by:
@@
III Transformers for Nonlinear Compensation
III-A Model Architecture
In this work, we use the above Transformer encoder architecture to estimate the nonlinear distortions in the received symbols. The overall design for the nonlinear channel equalization is depicted in Fig. 3. Due to the channel memory and pattern-dependency over the transmitted sequence, it is required to process several symbols before and after each target symbol. The number of these extra symbols from each side of the target symbol is denoted by tap-size value $t$. The model uses received symbols from four components of the optical domain at the end of the DSP chain, $X_{I},X_{Q},Y_{I},Y_{Q}$. At first, an embedding generator module maps this sequence of received symbols into $R^{N\times d_{model}}$, where $N$ is the embedding sequence length. Next, a Transformer is used as the core for nonlinear channel equalization. The dashed block in Fig. 2 corresponds to the Transformer block that operates on the generated embeddings. The last module processes the Transformer’s output to generate the equalizer’s nonlinear distortion estimations for one polarization (or for both polarizations by slightly modifying this output module).
Figure 3: The overall architecture for the nonlinear channel equalization. From left to right: the CNN generates input embeddings which are processed by the Transformer to generate output representations. These outputs are then fed into an MLP to generate the estimated nonlinear distortions $E_{XI}$ and $E_{XQ}$.
@@ III-A Model Architecture
For processing a block of symbols for equalization, the training/inference samples are created by selecting a block of $b$ symbols. In order to maintain the quality of equalization at the beginning and end of the block, we append extra symbols to the beginning and end of the block; effectively expanding the selected symbols with extra $t$ symbols at each side. Hence, each input sequence length is equal to $b+2t$. This sequence of symbols passes through the embedding generator module which produces $b+\ell$ embeddings (the exact value of $\ell$ will be explained later). These embeddings are then fed into the Transformer where $N=b+\ell$ outputs are generated.
The output module (MLP in the Fig. 3) generates $b$ estimated nonlinear distortions which are used for computing the loss in the training stage. This loss is generated as the mean of squared differences between $b$ estimated nonlinear distortions and $b$ target ones. Through back-propagation, the optimizer uses this loss to modify the weights in the training stage. With this approach, training can be accelerated by a large factor since instead of training over one symbol at a time, the model is trained on $b$ symbols at each training iteration. In addition, in the inference (equalization) stage, the Transformer generates $b$ nonlinear distortions at once through operations that are highly parallelizable. Block processing reduces the computational complexity of inference which can be seen from the following equation where the computational complexity of the attention matrix is expressed as a function of block size $b$:
$RMPS_{att-blk}=\frac{2h(b+\ell)^{2}d_{k}+3h(b+\ell)^{2}}{b}.$ (6)
@@
where $h$ is the number of heads and RMPS stands for real multiplications per symbol. In analyzing the computational complexity, we only consider the multiplication operations and do not include the additions. Hence, RMPS is used as the complexity metric in this work. The first term in the above equation is the complexity of matrix multiplications and the second term is the complexity of division by $\sqrt{d_{k}}$ and softmax computation. In a vanilla Transformer, the size of $d_{K}$ and $d_{V}$ are chosen as $d_{model}/h$. In our studies, we also explore the configurations where $d_{K}$ and $d_{V}$ are equal to $d/h$ where $d$ may not be equal to $d_{model}$. This case is particularly of interest since for $d<d_{model}$, we can reduce the computational complexity. Note that, ReLU and leaky ReLU’s complexities are excluded in our analysis, since their hardware implementations are fairly simple and requires very little resources. Note that the reduction factor of $b$ also exists in the other layers including the FFNs and output layer. Assuming $\ell$ and $h$ as constants, for large $b$, the overall complexity in Eq. (6) is approximately $O(bd_{k})$ per symbol which is in fact linear in block size $b$.
III-B Embedding
@@
Our studies over the course of this research show that the Transformer which directly operates on input symbols cannot effectively learn the nonlinear interactions among the symbols. Hence, we need to create good representations, commonly known as embeddings, of the input components, $X_{I},X_{Q},Y_{I},Y_{Q}$, prior to feeding them to the Transformer. Here, we explore two main classes of ANN models for generating the embeddings, an MLP and a CNN. For the MLP architecture, we examine two configurations. In the first one, an MLP with only one linear layer is used. In the second one, we employ a multi-layer MLP comprising a linear layer, a Leaky ReLU nonlinear activation with the negative slope of $0.2$, and another linear layer. Using these MLPs, we project the input symbols ($R^{(2t+b)\times 4}$) into embeddings with dimension $R^{(\ell+b)\times d_{model}}$ where $b$ is the processing block size, $\ell=2t$ and $t$ is the tap size as defined before. Note that the MLPs are applied similarly for each symbol over the input sequence.
@@
We also explore the use of CNNs for creating embeddings since they have shown great potential for generating features. We use a CNN with one layer of one-dimensional convolutional layer with a kernel size of $k$, four input channels, $d_{model}$ output channels, and a stride of one, followed by a leaky ReLU nonlinear activation function with negative slope of $0.2$. The CNN generates embeddings of dimension $R^{(\ell+b)\times d_{model}}$ where $\ell=2t-k+1$. Note that we use $\ell$ to unify our math notation for representing computation complexities for Transformers with either MLP or CNN embeddings-generator modules. In other words, $\ell$ is equal to $2t$ for MLP generated embeddings and $12t-k+$ 1 for CNN ones.
III-C Mask